OpenStudy (anonymous):

for intermediate algebra how do I solve the equation. 4/x-3-1=x-1/x-3

OpenStudy (tkhunny):

Lets work on notation a little. 4/x-3 mean \(\dfrac{4}{x} - 3\) x-1/x-3 means \(x - \dfrac{1}{x} - 3\) If you mean otherwise, you will need additional grouping symbols. 4/(x-3) mean \(\dfrac{4}{x-3}\) (x-1)/(x-3) means \(\dfrac{x - 1}{x - 3}\) Is that more like what you intended to solve?

OpenStudy (denisaboichuk):

x−1x−3